# algorithmic

### v2

@@ Appendix C Toy Experiments
C.2 Experimental result of constrained sample generation.
@@ D.1 Experimental design
Implementation. In this experiment, we used a pretrained denoiser $\epsilon_{\theta}(\cdot)$, and trained the noise level correction network $r_{\theta}(\cdot)$ to enhance the denoising process. For unconstrained image generation experiments, we employed the pretrained $32\times 32$ denoising network from Song et al. (2021a) in DDIM-based experiments on the CIFAR-10 dataset, and the pretrained $32\times 32$ denoising network from Karras et al. (2022) in EDM-based experiments on CIFAR-10. For image restoration experiments, we utilized the pretrained $256\times 255$ denoising network from Dhariwal & Nichol (2021) for ImageNet experiments and the pretrained $256\times 255$ denoising network from Lugmayr et al. (2022) for CelebA-HQ experiments.
Network architecture. The noise correction network $r_{\theta}(\cdot)$ is designed to be significantly smaller than the denoiser $\epsilon_{\theta}(\cdot)$ , while still incorporating residual and attention blocks similar to those used in the original DDPM denoiser Ho et al. (2020). Table 7 outlines the architecture of the noise correction network for CIFAR-10, ImageNet, and CelebA-HQ datasets. For all datasets, we employ 2 residual blocks, 1 attention block, and 4 attention heads. The primary architectural difference lies in the varying feature sizes (input channels and input dimensions) generated by the denoiser’s encoder. For comparison, the ImageNet denoiser contains 9 attention blocks in the encoder and 13 in the decoder. Consequently, the noise correction network is approximately ten times smaller than the denoiser network. The last two rows of Table 7 provide a parameter count comparison between the noise correction network $r_{\theta}$ and the denoiser $\sigma_{\theta}$.
@@
Network Heperparameter CIFAR-10 ImageNet Celeba-HQ
@@ D.2 Time and Memory Cost
Inference time. The inference times for 10-step sample generation with a batch size of 1 are presented in Table 9. As shown, incorporating noise level correction adds only a modest increase in inference time, approximately $\approx 10\%$.
Table 9: Inference time of DDIM and DDIM-NLC.
Inference time CIFAR-10 ImageNet
DDIM 0.32 0.93
DDIM-NLC 0.38 1.19
D.3 Lookup table method for image restoration

### v3

@@ Appendix C Toy Experiments
C.2 Misalignment between noise level and true distance to the manifold
As discussed in section 3.1, diffusion-based sample generation methods, such as DDIM and DDPM, often suffer from cumulative errors introduced during the denoising process. These errors arise from imperfections in the denoiser and become particularly significant in the final steps. As a result, there can be a substantial mismatch between the true distance to the data manifold, $\text{dist}_{\mathcal{K}}(x_{t})$, and the estimated noise level, $\sqrt{n}\sigma_{t}$. To visualize this phenomenon, we conduct a toy experiment. Figure 7 shows the deviation between the true distance $\text{dist}_{\mathcal{K}}(\hat{x}_{t})$ and the estimated noise level $\sqrt{n}\sigma_{t}$ during a 10-step DDIM sampling process. As observed, the actual distance increasingly deviates from the predefined noise level, especially in the later steps. This result aligns with the noise estimation inaccuracy illustrated earlier in fig. 2.
Figure 7: Deviation between the true distance $\text{dist}_{\mathcal{K}}(\hat{x}_{t})$ and the estimated noise level $\sqrt{n}\sigma_{t}$ in DDIM sampling.
C.3 Results of constrained sample generation.
@@ D.1 Experimental design
Network Hyperparameters CIFAR-10 ImageNet Celeba-HQ
@@
Implementation. In this experiment, we used a pretrained denoiser $\epsilon_{\theta}(\cdot)$, and trained the noise level correction network $r_{\theta}(\cdot)$ to enhance the denoising process. For unconstrained image generation experiments, we employed the pretrained $32\times 32$ denoising network from (Song et al., 2021a) in DDIM-based experiments on the CIFAR-10 dataset, and the pretrained $32\times 32$ denoising network from (Karras et al., 2022) in EDM-based experiments on CIFAR-10. For image restoration experiments, we utilized the pretrained $256\times 255$ denoising network from (Dhariwal & Nichol, 2021) for ImageNet experiments and the pretrained $256\times 255$ denoising network from (Lugmayr et al., 2022) for CelebA-HQ experiments.
Network architecture. The noise correction network $r_{\theta}(\cdot)$ is designed to be significantly smaller than the denoiser $\epsilon_{\theta}(\cdot)$ , while still incorporating residual and attention blocks similar to those used in the original DDPM denoiser (Ho et al., 2020). Table 7 outlines the architecture of the noise correction network for CIFAR-10, ImageNet, and CelebA-HQ datasets. For all datasets, we employ 2 residual blocks, 1 attention block, and 4 attention heads. The primary architectural difference lies in the varying feature sizes (input channels and input dimensions) generated by the denoiser’s encoder. For comparison, the ImageNet denoiser contains 9 attention blocks in the encoder and 13 in the decoder. Consequently, the noise correction network is approximately ten times smaller than the denoiser network. The last two rows of Table 7 provide a parameter count comparison between the noise correction network $r_{\theta}$ and the denoiser $\sigma_{\theta}$.
@@ D.2 Time and Memory Cost
Table 9: Inference Time Overhead of the Proposed NLC Method.
$\epsilon$ model CIFAR10 - DDPM CIFAR10-EDM CelebA - DDNM ImageNet-DDNM
Baseline Inference (s) 0.035 0.022 0.049 0.102
+NLC (s) 0.037 0.023 0.050 0.104
Time Overhead Ratio(%) 5.7 4.3 2.0 2.0
Table 10: Inference Time Comparison Between the Baseline and the Proposed NLC Method. For the CIFAR-10 experiment, the baseline is DDPM and the proposed method is DDPM+NLC. For the ImageNet experiment, the baseline is DDNM and the proposed method is DDNM+NLC (inpainting experiments).
$\epsilon$ model CIFAR10-DDPM (Song et al., 2021a) ImageNet - DDNM (Dhariwal & Nichol, 2021)
baseline 0.32 0.93
baseline + NLC 0.34 0.95
@@
Inference time. The proposed noise level correction (NLC) method requires only one additional inference step for the NLC network $r_{\theta}(\cdot)$ per overall inference. This extra step is independent of the specific sampling algorithm, applicable to both unconstrained and constrained approaches, as shown in line 3 of algorithms 1, 2, 3 and 4. Table 9 summarizes the additional time cost in different settings, indicating that each inference with the proposed NLC incurs an overhead of less than 6%. Furthermore, Table 10 presents the inference times for a real 10-step sample generation with a batch size of 1 for CIFAR-10 unconstrained sample generation with DDPM and ImageNet constrained (inpainting) generation with DDNM, confirming that incorporating the noise level correction results in a modest increase of approximately 6% in inference time.
D.3 Ablation Study on the Scaling Factor $\lambda$
Table 11: FID score on the scaling factor $\lambda$ in training $r_{\theta}(\hat{x}_{t},\sigma_{t})$. Here, $\delta=0$ corresponds to the naive setting ($\lambda=1$). CIFAR-10 experiments use 100-step sampling with DDIM-NLC, and ImageNet inpainting experiments use 100-step sampling with DDNM-NLC.
$\delta$ 0 0.2 0.5 1.0
CIFAR-10 (DDIM-NLC) 3.52 3.23 3.12 3.14
ImageNet Inpainting (DDNM-NLC) 7.87 7.59 7.20 7.24
In the training objective for the NLC network $r_{\theta}(\cdot)$, we introduce the scaling factor $\lambda$ to adjust the input-output range. Specifically, we use the predicted noise level $\sqrt{n}\sigma_{t}\bigl{[}1+r_{\theta}(\hat{x}_{t},\sigma_{t})\bigr{]}$ to estimate the true noise level (i.e., the distance from the noisy sample to the clean sample) $\|x_{t}-x_{0}\|$. This adjustment addresses the mismatch between the noise scales in the forward (diffusion) and reverse (denoising) processes. Noisy samples are generated as $\hat{x}_{t}=x_{0}+\sigma_{t}\lambda\epsilon_{t},$ where $\epsilon_{t}\sim\mathcal{N}(0,I)$ and typically $\|\epsilon_{t}\|\approx\sqrt{n}$. Without scaling (i.e., setting $\lambda\equiv 1$), the true noise level becomes concentrated around $\sigma_{t}\sqrt{n}$, leading to $r_{\theta}(\hat{x}_{t},\sigma_{t})\approx 0$. However, the denoiser network $\epsilon_{\theta}(x_{t},\sigma_{t})$ does not necessarily predict a noise level with such a concentrated norm. This norm mismatch, also noted in (Permenter & Yuan, 2024), motivates our introduction of $\lambda$.
By allowing $\lambda$ to vary, e.g., $\lambda\sim\mathcal{U}(1-\delta,1+\delta)$, the true noise level becomes approximately $\lambda\sigma_{t}\sqrt{n}$ and $r_{\theta}(\hat{x}_{t},\sigma_{t})$ reflects the variation in $\lambda$ (i.e., roughly in $[-\delta,\delta]$). To validate this design, we conduct an ablation study comparing different values of $\delta$ using 100-step sampling in terms of FID score, as shown in table 11. For CIFAR-10, the DDIM-NLC method is used for unconstrained sampling, while for ImageNet inpainting experiments the DDNM-NLC method is employed. As can be seen, $\delta=0.5$ (the standard setting of our experiment) outperforms the naive setting of $\delta=0$.
D.4 Lookup table method for image restoration
@@
D.5 Additional experiments on integrating NLC with DPM
Table 13: FID scores on CIFAR-10 using DPM sampling with and without NLC, across varying numbers of function evaluations (NFE).
[…]
DPM 5.28 3.43 3.02 2.85 2.78
DPM-NLC 4.83 3.22 2.97 2.81 2.77
As described in algorithm 5, we integrate the proposed Noise Level Correction (NLC) method into the widely used DPM-Solver sampling algorithm. To evaluate the effectiveness of this integration, we conduct experiments on the CIFAR-10 image generation task. Table 13 presents the FID scores for DPM with and without NLC under different numbers of function evaluations (NFE). As shown, incorporating NLC consistently improves sample quality across all settings. For instance, with NFE = 12, NLC reduces the FID score from 5.28 to 4.83, corresponding to a 9% improvement. These results demonstrate that NLC can effectively enhance the performance of DPM-based sampling methods.